Water from a dam falls on a turbine 30 m below at the rate of 60 m³/min. After





leaving the turbine wheel the water has a velocity of 9 m/s. What is the maximum





efficiency of the turbine? What is the maximum power developed?

1
Expert's answer
2021-12-16T14:39:17-0500

The maximum efficiency of the turbine


"\eta=\frac{W}{W_{total}}=\frac{\rho Qtgh-\rho Qtv^2/2}{\rho Qtgh}=1-\frac{v^2}{2gh}=1-\frac{9^2}{2\cdot9.8\cdot30}=0.86"


The maximum power developed


"P=\eta \rho Qt gh/t=\eta \rho Q gh=0.86\cdot1000\cdot1\cdot9.8\cdot30=252\ 840\ (W)"